# The Effectiveness of Large Language Models in Transforming Unstructured Text to Standardized Formats

### v1

@@ 3.2 Dataset
We assembled the dataset for this research from sources that provide recipes in both standard and Cooklang formats. The primary sources include the Cooklang documentation and associated GitHub repositories Cooklang (2024b, c); Cooklang Community (2024). We merged the recipes from these sources to create a comprehensive dataset, which we now make publicly available in our GitHub repository. GitHub repository [^2] https://github.com/williambrach/llm-text2cooklang/tree/main/data. The dataset contains 32 distinct recipe samples, representing a diverse range of recipe categories, including baking, breakfast, dinners, lunches, and soups. This diversity ensures a broad representation of culinary styles and techniques, enabling a more robust evaluation of the models. We enhanced the dataset by extracting ingredients from the Cooklang Specification using the Cooklang CLI CookLang (2024). We added these extracted ingredients as an additional feature to the dataset, enriching the information available for analysis and prompt optimization.
@@ 3.4 Hardware and Deployment Setup
We deployed all open-source models (Llama3.1:8b and LLama3.1:70b) locally using ollama Ollama (2024), a framework that runs and manages local deployment of LLMs. Our local setup included two NVIDIA GeForce RTX 4090 graphics cards, which provided the necessary computational power for efficient model inference. We accessed OpenAI GPT models through the OpenAI API.
@@ 4 Results
This section analyzes the performance of four Large Language Models (LLMs): GPT-4o, GPT-4o-mini, Llama3.1:70b, and Llama3.1:8b. The analysis evaluates each model using three traditional metrics: Word Error Rate (WER), ROUGE-L score, and Translation Edit Rate (TER), grouping all results by their mean values. Table 1 shows each model’s performance metrics:
@@
GPT-4o 0.1996 0.9280 0.8619
GPT-4o-mini 0.2907 0.8689 1.4088
Llama3.1:70b 0.9803 0.5453 4.7859
Llama3.1:8b 1.3513 0.4526 6.4910
@@
As shown in Figure 2, the GPT-4o model displays superior performance across all metrics in converting inputs to the Cooklang format. The model achieved the highest ROUGE-L score (0.9280), indicating an overlap between the generated Cooklang output and the reference text at the n-gram level. Furthermore, the model had the lowest Token Error Rate (TER) of 0.8619, indicating minimal hallucination or generation of content not grounded in the original text. Additionally, it showed the lowest Word Error Rate (WER) of 0.1996, implying high accuracy in maintaining the correct word sequence with minimal insertions, substitutions, or deletions when converting to Cooklang format. These results suggest that GPT-4o could be particularly well-suited for applications requiring accurate transformation between different text formats and structures, as its strong performance in maintaining semantic content while adapting syntactic structure indicates robust capabilities in text transformation tasks. GPT-4o-mini ranks second in converting plain text to Cooklang, demonstrating consistent performance, but falls behind GPT-4o. Its Word Error Rate (WER) of 0.2907 shows more errors in word sequence maintenance. The lower ROUGE-L score of 0.8689 reveals reduced overlap between the generated Cooklang and reference text. A Token Error Rate (TER) of 1.4088 indicates that GPT-4o-mini generates more content without grounding in the original text. Results for the family of Llama3.1 demonstrate models exhibit considerably reduced performance in Cooklang conversion relative to the GPT models. The 70b version exhibits superior performance compared to the 8b version across all metrics, suggesting that an increased model size contributes to enhanced performance. However, both models demonstrate a notable deficit in comparison to the GPT models. The WER values for the Llama models exceed 0.98, indicating a significant prevalence of errors in word sequence. The ROUGE-L scores are below 0.55, suggesting a notable discrepancy between the generated text and the reference text. Additionally, the TER values exceed 4.78, indicating a considerable tendency for hallucination or the generation of content that is not directly relevant to the source text.
@@ 4.1 Impact of Prompt Techniques on Model Performance
The efficacy of each prompt technique varies considerably across different models. Larger, more advanced models (GPT-4o and GPT-4o-mini) demonstrate superior overall performance and less variability between prompt techniques in comparison to smaller models (Llama3.1:70b and Llama3.1:8b). It is noteworthy that while Few-Shot consistently ranks first, the relative performance of MIPROv2 and Zero-Shot varies depending on the model and metric. Larger models demonstrate the most optimal overall performance across all techniques, with Few-Shot attaining a WER as low as 0.073 for GPT-4o. In contrast, Llama3 3.1 models exhibit elevated error rates and diminished ROUGE-L scores, with Zero-Shot outperforming MIPROv2, particularly for the 8b variant.
@@
GPT-4o (MIPROv2) 0.2221 0.9384 0.9950
GPT-4o (Few-Shot) 0.0730 0.9722 0.2096
GPT-4o (Zero-Shot) 0.3038 0.8733 1.3811
GPT-4o-mini (MIPROv2) 0.1962 0.8948 1.1047
GPT-4o-mini (Few-Shot) 0.0865 0.9646 0.2509
GPT-4o-mini (Zero-Shot) 0.6208 0.7387 2.9723
Llama3.1:70b (MIPROv2) 1.2060 0.4100 5.9615
Llama3.1:70b (Few-Shot) 0.8692 0.6170 4.4031
Llama3.1:70b (Zero-Shot) 0.8656 0.6091 3.9931
Llama3.1:8b (MIPROv2) 1.9504 0.2806 9.2570
Llama3.1:8b (Few-Shot) 1.1835 0.4898 5.8003
Llama3.1:8b (Zero-Shot) 0.9200 0.5874 4.4156
@@ 4.2 Effect of Cooking Language Specification and Ingredient Exclusion
We also evaluated if inserting Cooklang language specifications into prompts or if presenting a list of ingredients could yield a significant difference. Our findings indicate considerable discrepancies in model performance contingent on the utilization of Cooklang and the exclusion of ingredient information. Examined four models under different configurations: with/without Cooklang specification and with/without recipe ingredients. The results, presented in Table 3, demonstrate that these factors influence model performance across evaluation metrics.
@@
GPT-4o False False 0.3370 0.8839 1.4448
GPT-4o False True 0.1596 0.9392 0.6676
GPT-4o True True 0.1023 0.9608 0.4733
GPT-4o-mini False False 0.3924 0.8461 1.5210
GPT-4o-mini False True 0.2768 0.8800 1.2981
GPT-4o-mini True True 0.2075 0.8770 1.4443
Llama3.1:70b False False 1.1409 0.4483 5.3866
Llama3.1:70b False True 1.2036 0.4926 5.8035
Llama3.1:70b True True 0.5964 0.6951 3.1677
Llama3.1:8b False False 1.6832 0.3529 8.1012
Llama3.1:8b False True 1.2686 0.4796 5.7104
Llama3.1:8b False True 1.1021 0.5252 5.6614
@@ 4.3 Ingredient Identification
Our evaluation metric assessed how accurately LLMs identify all ingredients in recipes. Figure 6 illustrates the identification capabilities and shows performance variations across four tested models. GPT-4o achieved the highest score of 0.61 on the ingredient identification metric (0-1 range), while GPT-4o-mini scored 0.56. The LLama models performed significantly worse, with Llama3.1:70b scoring 0.24 and Llama3.1:8b reaching only 0.10. This gap reveals a significant difference between GPT and Llama models in ingredient identification tasks. The top GPT model outperforms the Llama model by approximately 2.5 times. Within the Llama family, increasing the model size from 8b to 70b parameters improved performance by 2.4 times, highlighting how the model scale affects this task. It is important to note that this evaluation methodology relies solely on exact string matching for ingredient identification without accounting for spelling variations or minor textual discrepancies. This strict matching criterion likely underestimates the models’ true ingredient identification capabilities, as it penalizes semantically correct but textually inexact matches. Future work could enhance this metric by incorporating fuzzy matching or semantic similarity measures to provide a more comprehensive assessment of ingredient identification performance. Figure 7 provides a deeper analysis, examining both missing ingredients (false negatives) and incorrectly added ingredients (false positives) across models.
@@
Figure 7: Comparison of models
@@ 4.4 Accuracy in Unit and Amount Identification
Table 4 shows LLM performance metrics across configurations and prompt techniques, focusing on unit and amount identification accuracy in domain-specific contexts. We evaluated the models using MIPROv2, few-shot, and zero-shot prompting techniques. GPT-4o, with few-shot prompting, achieved the highest scores, reaching 0.5729 and 0.5833 for unit and amount identification. Few-Shot and MIPROv2 techniques outperformed Zero-Shot across most models, highlighting effective prompting’s impact. GPT-4o and GPT-4o-mini surpassed the Llama models in performance, while the 70b Llama version exceeded its 8b variant. Zero-Shot yielded the lowest results, demonstrating proper prompting’s necessity. Llama models, particularly the 8b version, struggled compared to GPT models. Notably, Llama3.1:8b showed slightly better performance with Zero-Shot for amount detection. These results demonstrate how model size and prompting techniques affect specialized task performance while indicating room for improvement in domain applications. Similar to ingredient identification, in our metrics that evaluate LLMs across Unit and Amount Identification, the metric relies solely on exact string matching for ingredient identification without accounting for spelling variations or minor textual discrepancies.
@@
GPT-4o MIPROv2 0.5313 0.5625
GPT-4o Few-Shot 0.5729 0.5833
GPT-4o Zero-Shot 0.2292 0.2396
GPT-4o-mini MIPROv2 0.5234 0.5469
GPT-4o-mini Few-Shot 0.4688 0.4479
GPT-4o-mini Zero-Shot 0.1146 0.1563
Llama3.1:70b MIPROv2 0.1771 0.1667
Llama3.1:70b Few-Shot 0.1979 0.2083
Llama3.1:70b Zero-Shot 0.1042 0.1146
Llama3.1:8b MIPROv2 0.0104 0.0000
Llama3.1:8b Few-Shot 0.1146 0.1146
Llama3.1:8b Zero-Shot 0.1146 0.1250
@@ 5 Discussion
Our analysis of models, prompt techniques, and configurations demonstrates a clear pattern in the optimization of performance for both recipe processing and Cooklang conversion tasks. The results consistently indicate a hierarchy of performance, with GPT-4o exhibiting the highest performance, followed by GPT-4o-mini $>$ Llama 3.1:70b $>$ Llama 3.1:8b.
@@
  • Prompt Technique: The Few-Shot prompting technique demonstrated superior performance across all models. For GPT-4o specifically, Few-Shot achieved optimal results with WER=0.0730, ROUGE-L=0.9722, and TER=0.2096. MIPROv2 showed strong performance across most models, with Llama3.1:70b being the exception where both Zero-Shot and Few-Shot outperformed MIPROv2. The overall ranking of prompt techniques from our analysis: Few-Shot $>$ MIPROv2 $>$ Zero-Shot.
  • Cooklang and Ingredient Information: The implementation of the Cooklang format and ingredient information resulted in enhanced recipe parsing and generation, leading to improved WER, ROUGE-L, and TER scores. However, GPT models exhibited a tendency to include an excess of ingredients, occasionally introducing superfluous elements, as shown in Figure 7. In contrast, Llama models demonstrated a reduced proclivity to incorporate unnecessary ingredients but were more susceptible to the omission of essential components, potentially impacting the completeness and accuracy of the recipes.
@@
The results indicate that the optimal configuration for achieving optimal performance in recipe processing and Cooklang conversion tasks is the use of GPT-4o with Few-Shot prompting, which incorporates both Cooklang specifications and ingredient information. While GPT-4o-mini and larger Llama models (70b) demonstrated improvements with analogous configurations, the performance differential between GPT-4o and alternative models was considerable. However, the higher cost of using GPT-4o compared to smaller models presents a notable drawback, potentially limiting its practical implementation in resource-constrained scenarios where GPT-4o-mini could be more suitable with minimal performance reduction.
@@ 5.1 Open source model size comparison
The comparison between Llama3.1:8b and Llama3.1:70b reveals intriguing performance characteristics across all metrics. As anticipated, the larger 70b parameter model generally outperformed its 8b counterpart, consistent with the expectation that increased model size correlates with enhanced capabilities. However, a noteworthy observation emerges from the data shown in Figure 2: the upper-performance bounds of the 8b model frequently approach or overlap with those of the 70b model. This phenomenon is particularly evident in the WER and TER metrics, where the top quartile of Llama3.1:8b’s performance distribution nearly intersects with the median performance of Llama3.1:70b. The observed performance overlap indicates that, while the 70b model demonstrates more consistent high-level performance, the 8b model exhibits the potential for comparable results in optimal scenarios. This finding has significant implications for model selection and optimization strategies. The demonstrated capacity of the smaller model to achieve performance levels similar to its larger counterpart in certain instances opens avenues for targeted fine-tuning. By optimizing the 8b model for specific use cases, it may be possible to narrow or even eliminate the performance gap with the 70b model in particular applications. This potential for high performance in a smaller model through fine-tuning is particularly attractive from both practical and resource-efficiency perspectives.

### v2

@@ 3.2 Dataset
We assembled the dataset for this research from sources that provide recipes in both standard and Cooklang formats. The primary sources include the Cooklang documentation and associated GitHub repositories Cooklang (2024b, c); Cooklang Community (2024). We merged the recipes from these sources to create a comprehensive dataset, which we now make publicly available in our GitHub repository [^4] https://github.com/williambrach/llm-text2cooklang/tree/main/data. The dataset contains 1098 recipe samples, representing a range of recipe categories, including baking, breakfast, dinners, lunches, and soups. This diversity ensures a broad representation of culinary styles and techniques, enabling a more robust evaluation of the models. We enhanced the dataset by extracting ingredients from the Cooklang Specification using the Cooklang CLI CookLang (2024). We added these extracted ingredients as an additional feature to the dataset, enriching the information available for analysis and prompt optimization.
@@ 3.4 Hardware and Deployment Setup
We deployed all open-source models (Llama3.1:8b and LLama3.3:70b) locally using ollama Ollama (2024), a framework that runs and manages local deployment of LLMs. Our local setup included two NVIDIA GeForce RTX 4090 graphics cards, which provided the necessary computational power for efficient model inference. We accessed OpenAI GPT models through the OpenAI API.
3.5 Computational Overhead
Our methodology’s computational requirements can be analyzed across three phases:
  • Training: The DSPy framework introduces upfront computational overhead during training as it runs multiple prompt variations to select optimal configurations. This is a one-time cost; once optimal programs are identified, they can be reused during deployment.
  • Inference: The computational overhead during inference is negligible, primarily affecting the number of input tokens processed by the LLM. Since input token processing is less computationally expensive than output token generation, this represents a minor overhead compared to base LLM inference costs.
  • Implementation: The DSPy framework has minimal implementation overhead, efficiently constructing prompts without introducing significant processing loops or memory constraints. The memory footprint is minimal, consisting primarily of Python objects composed of strings.
In summary, our approach introduces a one-time training cost but maintains efficient performance during deployment with minimal additional computational requirements compared to standard LLM operation.
@@ 4 Results
This section analyzes the performance of four Large Language Models (LLMs): GPT-4o, GPT-4o-mini,LLama3.3:70b, and Llama3.1:8b. The analysis evaluates each model using three traditional metrics: Word Error Rate (WER), ROUGE-L score, and Translation Edit Rate (TER), grouping all results by their mean values. Table 1 shows each model’s performance metrics:
@@
GPT-4o 0.3893 0.8128 1.5255
GPT-4o-mini 0.4808 0.7550 2.0381
Llama3.3:70b 0.5510 0.7382 2.2802
Llama3.1:8b 0.5921 0.6911 2.6088
@@
As shown in Figure 2, the GPT-4o model displays superior performance across all metrics in converting inputs to the Cooklang format. The model achieved the highest ROUGE-L score (0.8128), indicating an overlap between the generated Cooklang output and the reference text at the n-gram level. Furthermore, the model had the lowest Token Error Rate (TER) of 1.5255, indicating minimal hallucination or generation of content not grounded in the original text. Additionally, it showed the lowest Word Error Rate (WER) of 0.3893, implying high accuracy in maintaining the correct word sequence with minimal insertions, substitutions, or deletions when converting to Cooklang format. These results suggest that GPT-4o could be particularly well-suited for applications requiring accurate transformation between different text formats and structures, as its strong performance in maintaining semantic content while adapting syntactic structure indicates robust capabilities in text transformation tasks. GPT-4o-mini ranks second in converting plain text to Cooklang, demonstrating consistent performance, but falls behind GPT-4o. Its Word Error Rate (WER) of 0.4808 shows more errors in word sequence maintenance. The lower ROUGE-L score of 0.7550 reveals reduced overlap between the generated Cooklang and reference text. A Token Error Rate (TER) of 2.0381 indicates that GPT-4o-mini generates more content without grounding in the original text. Results for the family of Llama3.1 demonstrate that models exhibit considerably reduced performance in Cooklang conversion relative to the GPT models. The 70b version exhibits superior performance compared to the 8b version across all metrics, suggesting that an increased model size contributes to enhanced performance. However, both models demonstrate a notable deficit in comparison to the GPT models. The WER values for the Llama models exceed 0.55, indicating a significant prevalence of errors in word sequence. The ROUGE-L scores are below 0.73, suggesting a similar performance as GPT-4o-mini in the generated text and the reference text. Additionally, the TER values exceed 2.28, indicating a possible tendency for hallucination or the generation of content that is not directly relevant to the source text.
@@ 4.1 Impact of Prompt Techniques on Model Performance
The efficacy of each prompt technique varies considerably across different models. Larger, more advanced models (GPT-4o and GPT-4o-mini) demonstrate superior overall performance and less variability between prompt techniques in comparison to smaller models (LLama3.3:70b and Llama3.1:8b). It is noteworthy that while Few-Shot consistently ranks first, but MIPROv2 optimized prompts are closely following. Zero-shot prompts show a consistent gap to other techniques for each model. Larger models demonstrate the most optimal overall performance across all techniques, with Few-Shot attaining a WER as low as 0.3509 for GPT-4o. In contrast, Llama3 3.1 models exhibit elevated error rates and lower ROUGE-L scores.
@@
GPT-4o (MIPROv2) 0.3637 0.8171 1.4566
GPT-4o (Few-Shot) 0.3509 0.8209 1.4467
GPT-4o (Zero-Shot) 0.4533 0.8005 1.6733
GPT-4o-mini (MIPROv2) 0.3664 0.8008 1.5024
GPT-4o-mini (Few-Shot) 0.3819 0.7928 1.5820
GPT-4o-mini (Zero-Shot) 0.6943 0.6715 3.0299
Llama3.3:70b (MIPROv2) 0.4037 0.8023 1.5364
Llama3.3:70b (Few-Shot) 0.3747 0.8270 1.4633
Llama3.3:70b (Zero-Shot) 0.8746 0.5854 3.8408
Llama3.1:8b (MIPROv2) 0.4869 0.7371 2.1442
Llama3.1:8b (Few-Shot) 0.4172 0.7837 1.8714
Llama3.1:8b (Zero-Shot) 0.8721 0.5526 3.8108
@@ 4.2 Effect of Cooking Language Specification and Ingredient Exclusion
We also evaluated whether inserting Cooklang language specifications into prompts or presenting a list of ingredients could yield a significant difference. Our findings indicate considerable discrepancies in model performance contingent on the utilization of Cooklang and the exclusion of ingredient information. Examined four models under different configurations: with/without Cooklang specification and with/without recipe ingredients. The results, presented in Table 3, demonstrate that these factors influence model performance across evaluation metrics.
@@
GPT-4o False False 0.4292 0.7585 1.8141
GPT-4o False True 0.3825 0.8350 1.4019
GPT-4o True True 0.3563 0.8449 1.3605
GPT-4o-mini False False 0.5429 0.7053 2.242
GPT-4o-mini False True 0.4970 0.7680 1.9757
GPT-4o-mini True True 0.4026 0.7919 1.8965
Llama3.3:70b False False 0.5972 0.6937 2.4857
Llama3.3:70b False True 0.6058 0.7329 2.3081
Llama3.3:70b True True 0.4501 0.7880 2.0466
Llama3.1:8b False False 0.5766 0.6863 2.4011
Llama3.1:8b False True 0.6127 0.7201 2.4220
Llama3.1:8b True True 0.5869 0.6669 3.0033
@@ 4.3 Ingredient Identification
Our evaluation metric assessed how accurately LLMs identify all ingredients in recipes. Figure 6 illustrates the identification capabilities and shows performance variations across four tested models. GPT-4o achieved the highest score of 0.33 on the ingredient identification metric (0-1 range), while GPT-4o-mini scored 0.296. The larger LLama3.3:70b closely followed GPT-4o-mini with 0.2589 (difference 0.0371). The smaller LLama3.1:8b performed worse and reaching only 0.216. This gap reveals a notable difference between GPT and Llama models in ingredient identification tasks. It is important to note that this evaluation methodology relies solely on exact string matching for ingredient identification without accounting for spelling variations or minor textual discrepancies. This strict matching criterion likely underestimates the models’ true ingredient identification capabilities, as it penalizes semantically correct but textually inexact matches. Future work could enhance this metric by incorporating fuzzy matching or semantic similarity measures to provide a more comprehensive assessment of ingredient identification performance. Figure 7 provides a deeper analysis, examining both missing ingredients (false negatives) and incorrectly added ingredients (false positives) across models.
@@
Figure 7: Comparison of inputting ingredients into model prompt
@@ 4.4 Accuracy in Unit and Amount Identification
Table 4 shows LLM performance metrics across configurations and prompt techniques, focusing on unit and amount identification accuracy in domain-specific contexts. We evaluated the models using MIPROv2, few-shot, and zero-shot prompting techniques. GPT-4o, with MIPROv2 optimized prompt, achieved the highest scores, reaching 0.3017 and 0.2900 for unit and amount identification. Few-Shot and MIPROv2 techniques outperformed Zero-Shot across most models, highlighting the impact of effective prompting. GPT-4o and GPT-4o-mini surpassed the Llama models in performance, while the 70b Llama version exceeded its 8b variant but, this metric was the hardest one for LLama3.3:70 with the largest gap it scored against GPT-4o-mini. Zero-shot yielded the lowest results, demonstrating the necessity of proper prompting. Llama models, particularly the 8b version, struggled compared to GPT models. Notably, Llama3.1:8b showed slightly better performance with Zero-Shot for amount detection. These results demonstrate how model size and prompting techniques affect specialized task performance while indicating room for improvement in domain applications. Similar to ingredient identification, our metrics that evaluate LLMs across Unit and Amount Identification rely solely on exact string matching for ingredient identification without accounting for spelling variations or minor textual discrepancies.
@@
GPT-4o MIPROv2 0.3017 0.29
GPT-4o Few-Shot 0.2433 0.2533
GPT-4o Zero-Shot 0.15 0.14
GPT-4o-mini MIPROv2 0.2033 0.2050
GPT-4o-mini Few-Shot 0.2250 0.2350
GPT-4o-mini Zero-Shot 0.0800 0.1317
Llama3.3:70b MIPROv2 0.0983 0.1183
Llama3.3:70b Few-Shot 0.1600 0.1900
Llama3.3:70b Zero-Shot 0.1033 0.1550
Llama3.1:8b MIPROv2 0.1250 0.1533
Llama3.1:8b Few-Shot 0.1433 0.1750
Llama3.1:8b Zero-Shot 0.0817 0.1033
@@ 5 Discussion
Our analysis of models, prompt techniques, and configurations demonstrates a clear pattern in the optimization of performance for both recipe processing and Cooklang conversion tasks. The results consistently indicate a hierarchy of performance, with GPT-4o exhibiting the highest performance, followed by GPT-4o-mini $>$ Llama 3.3:70b $>$ Llama 3.1:8b. We included GPT-4o, GPT-4o-mini, LLama3.3:70b, and Llama3.1:8b to represent different points in the spectrum of model scale and availability. Our results do show that GPT-4o performed best on our specific task, but this finding was reported as an empirical observation rather than advocacy. Results could be skewed towards the OpenAI family models because the compared models from the Llama family were significantly smaller.
@@
  • Prompt Technique: The Few-Shot prompting technique demonstrated best performance across all models. For GPT-4o specifically, Few-Shot achieved optimal results with WER=0.3509, TER=1.4467 but was outperformed in ROUGE-L metrics by Llama3.3:70b where llama model achvied score 0.8270 and GPT-4o achvied 0.8209. MIPROv2 showed strong performance across most models. The overall ranking of prompt techniques from our analysis: Few-Shot $>$ MIPROv2 $>$ Zero-Shot.
  • Cooklang and Ingredient Information: The implementation of the Cooklang format and ingredient information resulted in enhanced recipe parsing and generation, leading to improved WER, ROUGE-L, and TER scores. However, GPT models exhibited a tendency to include an excess of ingredients, occasionally introducing superfluous elements, as shown in Figure 7. In contrast, Smaller llama model demonstrated a reduced proclivity to incorporate unnecessary ingredients but were more susceptible to the omission of essential components, potentially impacting the completeness and accuracy of the recipes.
@@
The results indicate that the optimal configuration for achieving optimal performance in recipe processing and Cooklang conversion tasks is the use of GPT-4o with Few-Shot prompting, which incorporates both Cooklang specifications and ingredient information, and GPT-4o-mini and the larger Llama model (70b) demonstrated improvements with analogous configurations and closely following GPT-4o in results. However, the higher cost of using GPT-4o compared to smaller models presents a notable drawback, potentially limiting its practical implementation in resource-constrained scenarios where GPT-4o-mini or Llama3.3:70b could be more suitable with minimal performance reduction.
@@ 5.1 Open source model size comparison
The comparison between Llama3.1:8b and Llama3.3:70b reveals significant performance differences across all metrics. As anticipated, the larger 70b parameter model outperformed its 8b counterpart, consistent with the expectation that increased model size correlates with enhanced capabilities. The data presented in Figure 2 and Table 1 demonstrates a clear performance hierarchy: the 70b model achieves a WER of 0.5510, ROUGE-L of 0.7382, and TER of 2.2802, compared to the 8b model’s WER of 0.5921, ROUGE-L of 0.6911, and TER of 2.6088. This substantial performance gap between the two open-source models indicates that the 8b model’s performance distribution does not significantly overlap with that of the 70b model. For scenarios where deployment constraints favor open-source alternatives, the substantial performance advantage of the 70b model over the 8b variant likely justifies the additional computational resources required, particularly for applications where accuracy is paramount.
5.2 Scalability Considerations
Our evaluation highlights critical scalability considerations for real-world implementation. GPT-4o demonstrated superior performance but introduced significant deployment challenges via API, including latency (average processing time of 3 seconds per recipe) and cost (approximately \$8.75 per 1,000 recipes with 1500 input/500 output tokens). GPT-4o-mini offers a more economical alternative at \$0.53 per 1,000 recipes with reduced latency (1 second per recipe), though both figures assume no parallel API calls. In contrast, locally deployed Llama3.1:8b processed 40 recipes per minute on consumer-grade hardware (single Nvidia RTX 4090), eliminating API costs and significantly reducing processing time, making it suitable for batch processing large recipe collections despite lower accuracy. For large-scale applications processing thousands of recipes (or other domain-specific text data) daily, the performance-cost trade-off becomes significant. Additional scalability challenges include handling diverse recipe formats and languages, managing concurrent requests, and integrating with existing culinary databases. These considerations will become increasingly important as this technology is deployed beyond experimental settings to production environments.